OpenStudy (anonymous):

The marked price of a car is 50% higher than the cost. If the car is sold at a 20% discount on the marked price, then the percentage profit is __%?

OpenStudy (phi):

any idea ?

OpenStudy (phi):

up 50% means the price is 1.5 times the original (100% + 50% = 1+.5=1.5) 20% discount means 100%-20% = 1-0.2 = 0.8 multiply by 0.8 so all together 1.5*0.8 = the final price subtract 1 to get the profit. change to a percent to get percent profit

OpenStudy (anonymous):

20%?

OpenStudy (phi):

yes

OpenStudy (anonymous):

THZ